Given the demand function Qd = 400 - Q2

Compute the price elasticity of demand when P= 10, and comment on the elasticity of the good.
1
Expert's answer
2020-11-23T10:22:04-0500

"\bold {Answer}"

"\eta_d =" "-0.667"


It is Inelastic since "\left| -0.667 \right | < 1"


"\bold {Solution}"

We are given "Q_d = 400 - P^2"


When "P = 10," "Q_d = 400 - (10)^2"

"= 400-100"

"= 300"



"\eta_d = \dfrac {dQ}{dP} × \dfrac {P_0}{Q_0}"


"\dfrac {dQ}{dP} = \dfrac {d}{dP} (400-P^2)"


"= -2P"


When "P = 10," "\dfrac {dQ}{dP} = -2(10)"


"= -20"


Therefore, "\eta_d = -20 × \dfrac {10}{300}"


"=- \dfrac {200}{300}"


"= -0.666666667"

"= \bold {-0.667}"


"\therefore \space" the elasticity of demand is "\bold {-0.667}" and it is "\bold {inelastic}."